OpenStudy (anonymous):

Find the center and radius of x^2 + y^2 + 2x - 10y + 10 = 0 a. center (1, - 5); r = 16 b. center (1, -5); r = 4 c. center (-1, 5); r = 16 d. center (-1, 5); r= 4

OpenStudy (mertsj):

x^2+2x+1 +y^2-10y+25 = -10+1+25 (x+1)^2 +(y-5)^2 = 16 Center (-1,5) r=4